A positive ion with charge 2e and mass 3.4 X 1027 kg passes through the velocity selector and into the detection chamber of a mass spectrometer as shown. In the velocity selector the electric field E-7000.0 V/m to the left and the magnetic field Bi 0.09 T out of the page. a) What is the speed v of the ion? (5 points) b) When the ion enters the upper detection chamber, it bends to the right with a radius of 10.0 mm. What is the magnitude and direction of B,? (5 points)

Please help and show work, thank you
Answer #1

a) since the charge particle is able to move in upper chamber, it means its path must be straight line, and this could be happen if lorentz force acting on charge particle in velocity selector would be zero

Lorentz forfe is given by

F = q ( E + (-Bv))

0 = E - Bv

v = E/B = 7000/0.09

v = 7.778*10^4 m/s

=============

As we know, radius executed by moving charge particle in magnetic field is given by

r = mv / qB

B = mv/qr

B = 3.4*10^-27* 7.778*10^4/( 2*1.6*10^-19* 10*10^-3)

B = 0.08264 T

===========

Comment in case any doubt.. Goodluck